OpenStudy (anonymous):

(t-5)^2=-28 solve using the square root property

OpenStudy (anonymous):

no real solutions

OpenStudy (anonymous):

but for complex numbers\[(t-5) = +- \sqrt{28} i \]\[t = 5 +- \sqrt{28}i\]

OpenStudy (anonymous):

Wouldn't it be t=5+-2i square root 7?

OpenStudy (anonymous):

yeh you could take the 4 out of the sqrt if you wanted to , i didnt really care that much about the little simplifications

OpenStudy (anonymous):

Ok. I have to take it down that far for the instructor. Thanks for the help!